# Fast data inversion for high-dimensional dynamical systems from noisy measurements

### v2

@@ Abstract
Bayesian prior,
latent factor models,
Gaussian processes,
expectation-maximization algorithm,
Kalman filter,
@@
, and
1 Introduction
Latent factor models of time-dependent data have wide applications, including modeling multiple time series in economics (Lam and Yao,, 2012) and estimating unobservable geophysical processes by ground deformation measurements (Segall and Matthews,, 1997). Though applications differ, the latent factor model can be broadly classified into two categories with either fixed or data-dependent factor loading matrices.
Basis functions, such as discrete Fourier basis and Green functions, are widely used for inverse estimation of experimental or field observations. Differential dynamic microscopy (Cerbino and Trappe,, 2008), for instance, is a physical approach for estimating the rheological properties of the materials by microscopy videos. The estimation corresponds to minimizing the temporal autocorrelation in the Fourier space by a latent factor model with the complex conjugate of the discrete Fourier basis (Gu et al., 2024a, ). Green’s functions, as another example, relate field observations to unobserved displacement at different spatial scales, such as cellular force estimation by traction force microscopy (Sabass et al.,, 2008) and slip estimation by geodetic data (Yabuki and Matsu’Ura,, 1992).
On the other hand, the factor loading matrix can be estimated by data. Probabilistic models were built to understand the underlying model assumptions made by these estimations. The loading matrix estimated by the principal component analysis (Berkooz et al.,, 1993), for instance, is shown to be the maximum marginal likelihood estimator of a latent factor model with random factors independently distributed as standard Gaussian distributions (Tipping and Bishop,, 1999). Various approaches extend the independent assumption for correlated data. For instance, dynamic mode decomposition is a popular approach of linearizing the one-step-ahead transition operator of nonlinear dynamical systems and reconstructing the dynamics by eigenpairs of the linear mapping matrix (Schmid,, 2010; Tu et al.,, 2014). In Gu et al., 2024b , the authors show the estimation of the dynamic mode decomposition is equivalent to the maximum likelihood estimator of the linear mapping matrix in a vector autoregressive model of noise-free observations. As another example, each latent factor is modeled by a Gaussian process in coregionalization models (Gelfand et al.,, 2004).
In this work, we develop a scalable and efficient approach for latent factor models with correlated factors and an orthogonal factor loading matrix either fixed or estimated by data. Our contributions are threefold. First, estimating a large number of correlation parameters in multivariate Gaussian processes is a fundamental challenge. When each latent factor process has distinct parameters, conventional strategies, such as posterior sampling or numerical optimization can be prohibitively slow for estimating a large number of parameters. We developed a new expectation-maximization (EM) algorithm. We surprisingly found that the estimation of the factor loading matrix, correlation and variance parameters, all have a closed-form expression in the algorithm. In particular, the correlation parameters of latent processes can be solved by cubic equations of order three. Furthermore, we construct an orthogonal latent factor loading matrix which enables us to filter each projected latent factor process separately without approximation, thereby avoiding large matrix inversion conventionally needed at each time in the Kalman filter (Kalman,, 1960). Second, we model latent factors by a multivariate Ornstein-Uhlenbeck process that contains distinct correlation and variance parameters. Our model extends the probabilistic model of the dynamic mode decomposition by including the noise model and utilizing a symmetric factor loading matrix for reducing the space in estimation. Third, we developed a new way to estimate the number of factors with either fixed or estimated factor loading matrix by matching the estimated noise variance with its measurement. Our study is motivated by fault slip estimation from geodetic measurements, but the new approach has broad applications for problems in science and engineering.
1.1 Motivating example: Fault slip estimation by geodetic data
Ground deformation observations, such as continuous Global Positioning System (GPS) observations (Segall and Matthews,, 1997) and interferometric synthetic-aperture radar interferograms (Anderson et al.,, 2019), have been widely used to quantify geological hazards, including volcanic eruptions and earthquakes. The goal of our application is to estimate fault slip that quantifies the relative movement velocities of two sides of the faults (Bürgmann,, 2018). Since slip cannot be directly observed, time-dependent ground deformation information from GPS data has been used for slip estimation (Bock et al.,, 1993).
[…]
@@ 1.1 Motivating example: Fault slip estimation by geodetic data
Suppose we collect GPS observations at $\tilde{k}$ locations, resulting in a $k=p_{y}\tilde{k}$ vector of observations $\mathbf{y}(t)=[\mathbf{y}(\mathbf{x}_{1},t),\mathbf{y}(\mathbf{x}_{2},t),\dots,\mathbf{y}(\mathbf{x}_{\tilde{k}},t)]^{T}$. In Segall and Matthews, (1997), the output vector is modeled below:
$\mathbf{y}(t)={\mathbf{G}}\mathbf{z}_{s}(t)+\bm{\epsilon}(t),$ (1)
@@
where $\mathbf{G}$ is a $k\times k^{\prime}$ matrix of discretized Green’s function, $\mathbf{z}_{s}(t)=[z_{s}(\bm{\xi}_{1},t),\dots,z_{s}(\bm{\xi}_{k^{\prime}},t)]^{T}$ is a $k^{\prime}$-vector of unobservable fault slip, with subscript ‘s’ meaning the slip, and the Gaussian noises follow $\bm{\epsilon}(t)\sim\mathcal{MN}(\mathbf{0},\sigma^{2}_{0}\mathbf{I}_{k})$ with variance $\sigma^{2}_{0}$. The $(i,h)$ block of $\mathbf{G}$ is $\mathbf{G}_{i,h}=\mathbf{G}(\mathbf{x}_{i},\bm{\xi}_{h})\Delta$ and $\Delta$ is the area size in discretization, for $i=1,...,\tilde{k}$ and $h=1,...,k^{\prime}$ with $\tilde{k}=100$ GPS stations and $k^{\prime}=1978$ discretization points used in estimating the slip propagation in Cascadia (Bartlow et al.,, 2011). We found that increasing the number of discretization points has almost no impact on estimates of the slip.
@@
Let $\mathbf{U}_{0}$ be $k\times d$ left singular vector matrix corresponding to the largest $d$ singular values in the singular value decomposition (SVD) of the Green’s function $\mathbf{G}\approx\mathbf{U}_{0}\mathbf{D}_{0}\mathbf{V}_{0}^{T}$, where $\mathbf{D}_{0}$ is a $d\times d$ diagonal matrix and $\mathbf{V}_{0}$ is a $k^{\prime}\times d$ matrix of $d$ right singular vectors. As the number of discretization points $k^{\prime}$ can be large, the latent slip vector is modeled by a linear combination of the latent processes (Segall and Matthews,, 1997):
$\mathbf{z}_{s}(t)={\mathbf{G}}^{T}\mathbf{U}_{0}\mathbf{\tilde{z}}(t),$ (2)
@@
with $\mathbf{\tilde{z}}(t)$ being a $d$-vector of latent processes. The latent process $\tilde{z}_{l}(t)$ is assumed to follow an integrated Brownian motion in Segall and Matthews, (1997). The Kalman filter is used for computing the posterior distribution of the latent process $\mathbf{\tilde{z}}(\cdot)$ and slip estimation. There are several restrictions in this approach. First, the latent process ${\tilde{z}}_{l}(\cdot)$ has the same covariance across $l=1,...,d$, yet the random latent factors can have different correlation length scales. Second, one needs to invert a $k\times k$ covariance matrix at each time in the Kalman filter, which is prohibitively slow when the number of GPS stations is moderately large. Third, selecting the number of latent factors is not discussed in this approach.
@@
The goal of this work is to develop a generally applicable approach and scalable algorithm for latent models for estimating the mean of the data and the posterior distribution of the latent variables. For estimating the slip rates in Cascadia between June and August in 2011, we found that the estimated slips had higher spatial correlation with independently detected tremor events not used in estimation. Furthermore, our method is $300-3000$ times faster than the current method in our real application. The high scalability of our method enables abundant data from GPS networks to be jointly used for hazard quantification.
The rest of the paper is organized as follows. In Section 2, we introduce a flexible latent factor model for high-dimensional dynamical systems with a fast EM algorithm for estimation. We show our model substantially accelerates the computation compared to other approaches in Section 3. Extensive experiments in Section 4 show high accuracy and computational scalability of our approach. In Section 5, we compare our approach with existing ways to estimate the slip migration in the Cascadia region by the GPS data and demonstrate the higher detection rate of tremor events from the independent source of seismic data not used in estimation. We conclude our study and outline future directions in Section 6.
@@ 2.1 A multivariate Ornstein-Uhlenbeck process with an orthogonal basis matrix
We first introduce our model and its properties. The derivations in this subsection are provided in Section S1 in the supplementary materials. We follow the assumption in (2), which indicates that the mean in (1) is approximated by ${\mathbf{G}}\mathbf{z}_{s}(t)\approx\mathbf{U}_{0}\mathbf{z}(t)$, where $\mathbf{z}(t)=\mathbf{D}^{2}_{0}\mathbf{\tilde{z}}(t)$ with the $l$th entry being $z_{l}(t)$, and there is no approximation when $k=d$. We model the $d$-dimensional latent factor process $\mathbf{z}(\cdot)$ by independent Ornstein-Uhlenbeck processes, where each process has its own correlation and variance parameters. Putting together, Equations (1)-(2) lead to
$\mathbf{y}(t)$ $=\mathbf{U}_{0}{\mathbf{z}}(t)+\bm{\epsilon}(t),$ (3)
${z}_{l}(t)$ $=\rho_{l}{z}_{l}(t-1)+w_{l}(t),$ (4)
where $\mathbf{U}_{0}$ is a $k\times d$ orthogonal factor loading matrix, either fixed or estimated, which relates the $d$-dimensional latent factors to $k$-dimensional measurements. Here, the parameter $\rho_{l}$ controls the temporal correlation length of the $l$th latent process, and $w_{l}(t)\sim\mathcal{N}(0,\sigma^{2}_{l})$ represents the innovation of the $l$th latent process for $t=2,...,n$, with the initial state being $z_{l}(1)\sim\mathcal{N}(0,\tau^{2}_{l})$ for $l=1,\dots,d$. We postpone the discussion of estimating $d$ in Section 2.3. Note that the matrix $\mathbf{G}$ in Equation (1) is typically neither symmetric nor orthogonal, whereas $\mathbf{U}_{0}$ is an orthogonal matrix, which brings substantial computational advantages, as will be elaborated soon. In this work, we discuss algorithms for both fixed and estimated $\mathbf{U}_{0}$, which are broadly applicable to many applications. Furthermore, we assume the mean to be zero for simplicity, though additional mean structures can be included in the model.
Lemma 2.1 below connects model in (3)-(4) to the continuous-time multivariate Ornstein-Uhlenbeck process (Gardiner,, 1985; Meucci,, 2009) with an orthogonal basis matrix.
@@
$d\mathbf{m}(t)=-\mathbf{U}_{0}\mathbf{D}\mathbf{U}^{T}_{0}\mathbf{m}(t)dt+\mathbf{U}_{0}\mathbf{\tilde{D}}d\mathbf{B}_{t},$ (5)
@@
We study the model in (3)-(4) because of its flexibility due to distinct parameters $(\rho_{l},\sigma_{l}^{2})$ for each latent factor $l$, for $l=1,...,d$ and the computational scalability when the number of latent factors, $d$, is large. The latent process $z_{l}(\cdot)$ can be either stationary or nonstationary, depending on the assumption of the initial state, stated in the Lemma 7 below.
@@
$\mbox{Cov}[z_{l}(t),z_{l}(t^{\prime})]=\rho_{l}^{t^{\prime}-t}\left\{\rho_{l}^{2(t-1)}\tau_{l}^{2}+\frac{1-\rho_{l}^{2(t-1)}}{(1-\rho_{l}^{2})}\sigma_{l}^{2}\right\},\quad\text{for }t^{\prime}\geq t.$ (6)
In particular, when the variance of initial state is $\tau^{2}_{l}=\frac{\sigma^{2}_{l}}{(1-\rho^{2}_{l})}$, $z_{l}(\cdot)$ is stationary with
$\mbox{Cov}[z_{l}(t),z_{l}(t^{\prime})]=\frac{\sigma^{2}_{l}}{1-\rho^{2}_{l}}\rho_{l}^{|t-t^{\prime}|}.$ (7)
In the rest of the article, we assume that each latent process is stationary by letting $\tau^{2}_{l}=\frac{\sigma^{2}_{l}}{1-\rho^{2}_{l}}$. Lemma 2.3 outlines key properties of the covariance and precision matrices, which enable us to modify Kalman filter (KF) (Kalman,, 1960) and Rauch–Tung–Striebel (RTS) smoother (Rauch et al.,, 1965) to compute several key quantities in our fast algorithm.
@@
Assume $\tau^{2}_{l}=\frac{\sigma^{2}_{l}}{1-\rho^{2}_{l}}$ and denote $\mathbf{\Sigma}_{l}=\frac{\sigma_{l}^{2}}{(1-\rho_{l}^{2})}\mathbf{R}_{l}$ as the covariance matrix of $\mathbf{z}_{l}=(z_{l}(1),\dots,z_{l}(n))^{T}$ defined in Equation (4). The $(t,t^{\prime})$ entry of $\mathbf{R}_{l}$ is $\rho_{l}^{|t-t^{\prime}|}$, for $t,t^{\prime}=1,2,\dots,n$. The inverse of $\mathbf{R}_{l}$ has a tri-diagonal structure, with diagonal entries of $\frac{1}{1-\rho_{l}^{2}}$ at the first and last positions, and $\frac{1+\rho_{l}^{2}}{1-\rho_{l}^{2}}$ at the remaining positions. The primary off-diagonal entries are $-\frac{\rho_{l}}{1-\rho_{l}^{2}}$. Additionally, the determinant of $\bm{\Sigma}_{l}$ follows $|\bm{\Sigma}_{l}|=\frac{\sigma^{2n}}{1-\rho_{l}^{2}}$.

### v3

@@ Abstract
Bayesian prior, latent factor models, Gaussian processes, expectation-maximization algorithm, Kalman filter
{AMS}
62-04, 62-08, 62C12, 62H11, 62H25, 62M20, 62M30, 62P35
1 Introduction
Latent factor models of time-dependent systems have wide applications, such as estimating unobservable geophysical processes [33], model calibration for computer simulation of multivariate outputs [19], and inferring multiple time series in economics [22]. Though applications differ, latent factor models can be broadly classified into two categories with either fixed or data-dependent factor loading matrices.
Basis functions, such as discrete Fourier basis and Green’s functions, are widely used for inverse estimation of experimental or field observations. Differential dynamic microscopy [9], for instance, is a physical approach for estimating the rheological properties of the materials by microscopy videos. The estimation corresponds to minimizing the temporal autocorrelation in the Fourier space by a latent factor model with the complex conjugate of the discrete Fourier basis [16]. Green’s functions, as another example, relate field observations to unobserved displacement at different spatial scales, such as cellular force estimation by traction force microscopy [31] and geologic slip estimation by geodetic data [44].
On the other hand, the factor loading matrix can be estimated by data. Probabilistic models were built to understand the underlying model assumptions made by these estimations. The loading matrix estimated by the principal component analysis [5], for instance, is shown to be the maximum marginal likelihood estimator of a latent factor model with random factors independently distributed as standard Gaussian distributions [38]. Various approaches extend the independent assumption for correlated data. In [17], for instance, the authors show the estimation of the dynamic mode decomposition is equivalent to the maximum likelihood estimator of the linear mapping matrix in a vector autoregressive model of noise-free observations. As another example, each latent factor is modeled by a Gaussian process in coregionalization models [13].
In this work, we develop a scalable and efficient approach for latent factor models with correlated factors and an orthogonal factor loading matrix, either fixed or estimated by data. Our contributions are threefold. First, estimating a large number of correlation parameters in multivariate Gaussian processes is a fundamental challenge. When each latent factor process has distinct parameters, conventional strategies, such as posterior sampling or numerical optimization can be prohibitively slow for estimating a large number of parameters. We developed a novel expectation-maximization (EM) algorithm for fast parameter estimation, which avoids expensive matrix inversion at each time in the Kalman filter required in the conventional EM algorithm for state space models [34] based on the orthogonal projection through the latent factor loading matrix. We surprisingly found that the estimation of the factor loading matrix, correlation, and variance parameters all have a closed-form expression in the algorithm. In particular, the correlation parameters of latent processes can be solved by cubic equations of order three. These key results are provided in Theorem 2.5. The log likelihood in the EM algorithm typically converges in less than 20 iterations in most of the scenarios studied in this work, an example of which is shown in Figure 4. The fast convergence makes our fast approach an almost exact solution for many problems. Second, we model latent factors by a multivariate Ornstein-Uhlenbeck process that contains distinct correlation and variance parameters, which provide a flexible way to capture dependence across output coordinates. Third, we develop a new way to estimate the number of factors with either a fixed or an estimated factor loading matrix by matching the estimated noise variance with its measurement.
Our approach has advantages over alternative approaches developed in computational mathematics, statistics, and geophysics communities. In computational mathematics, dynamic mode decomposition is a popular approach that linearizes the one-step-ahead transition operator of nonlinear dynamical systems to reconstruct the dynamics by the eigenpairs of the linear mapping matrix [32, 39], which produces a finite-dimensional approximation of the Koopman modes and eigenvalues [27]. Our model extends the probabilistic model of the dynamic mode decomposition by including the noise model and utilizing a symmetric factor loading matrix for reducing the space in estimation. Significant improvements against dynamic mode decomposition will be shown for noisy observations in Section 4.
In the statistics community, EM algorithm and Kalman filter have long been used for estimating parameters in the vector autoregressive models [34, 25], and they are implemented in [36]. Our approach contains three advantages against the conventional EM algorithm for autoregressive models. First, our approach enables both the estimated factor loading matrix and dimension reduction of latent factors, while the EM algorithm typically assumes that the dimension of the latent factor is the same as the number of time series. Second, assuming models with $d$ latent factor processes at $n$ time points, our new algorithm only requires $\mathcal{O}(nd)$ operations in Kalman filter without approximation due to the orthogonality of the latent factor loading matrix, while the conventional EM algorithm for vector autoregressive models requires $\mathcal{O}(nk^{3})$ operations for inverting a $k\times k$ one-step-ahead predictive covariance matrix of the observations at each time step in Kalman filter. Third, our model can be set to be either stationary or non-stationary, yet the estimation from the conventional EM algorithm cannot guarantee the stationarity [34].
In the geophysics community, the network inverse filter [33, 4] is widely used for estimating geologic slips from ground deformation measurements, such as the GPS time series. It assumes the vector autoregressive processes with the latent states following integrated Brownian motions, with shared parameters across different processes. Our method is both theoretically faster and more flexible with distinct correlation and variance parameters for each latent process, as will be shown in Section 4 and Section 5.
The rest of the paper is organized as follows. In Section 2, we introduce a motivating example for estimating geologic slip and introduce a flexible latent factor model for high-dimensional dynamical systems with a fast EM algorithm given in Theorem 2.5. We show our model substantially accelerates the computation compared to other approaches in Section 3. Extensive experiments in Section 4 show high accuracy and computational scalability of our approach. In Section 5, we compare our approach with existing ways to estimate the slip migration in the Cascadia region using the GPS data and demonstrate the higher detection rate of tremor events from an independent source of seismic data not used in estimation. We conclude our study and outline future directions in Section 6. The FMOU algorithm is coded in the FastGaSP package available on CRAN [15]. The data and code are made publicly available in GitHub: https://github.com/UncertaintyQuantification/FMOU/.
2.1 Motivating example: Geologic slip estimation by the ground deformation data
Our study is motivated by geologic slip estimation from geodetic measurements, but the new approach has broad applications for problems in science and engineering. Ground deformation observations, such as continuous Global Positioning System (GPS) observations [33] and interferometric synthetic-aperture radar interferograms [2], have been widely used to quantify geological hazards, including volcanic eruptions and earthquakes. The goal of our application is to estimate slip that quantifies the relative movement velocities of two sides of the faults [7]. Since slip cannot be directly observed, time-dependent ground deformation information from GPS data has been used for slip estimation [6].
@@ 2.1 Motivating example: Geologic slip estimation by the ground deformation data
Suppose we collect GPS observations at $\tilde{k}$ locations, resulting in a $k=p_{y}\tilde{k}$ vector of observations $\mathbf{y}(t)=[\mathbf{y}(\mathbf{x}_{1},t),\mathbf{y}(\mathbf{x}_{2},t),\dots,\mathbf{y}(\mathbf{x}_{\tilde{k}},t)]^{T}$. In [33], the output vector is modeled as:
(1) $\mathbf{y}(t)={\mathbf{G}}\mathbf{z}_{s}(t)+\bm{\epsilon}(t),$
@@
where $\mathbf{G}$ is a $k\times k^{\prime}$ matrix of discretized Green’s function, $\mathbf{z}_{s}(t)=[z_{s}(\bm{\xi}_{1},t),\dots,z_{s}(\bm{\xi}_{k^{\prime}},t)]^{T}$ is a $k^{\prime}$-vector of unobservable geologic slip, with subscript ‘s’ meaning the slip, and the Gaussian noises follow $\bm{\epsilon}(t)\sim\mathcal{MN}(\mathbf{0},\sigma^{2}_{0}\mathbf{I}_{k})$ with variance $\sigma^{2}_{0}$. The $(i,h)$ block of $\mathbf{G}$ is $\mathbf{G}_{i,h}=\mathbf{G}(\mathbf{x}_{i},\bm{\xi}_{h})\Delta$ and $\Delta$ is the area size in discretization, for $i=1,...,\tilde{k}$ and $h=1,...,k^{\prime}$ with $\tilde{k}=100$ GPS stations and $k^{\prime}=1978$ discretization points used in estimating the slip propagation in Cascadia [4]. We found that increasing the number of discretization points has almost no impact on estimates of the slip.
@@
Let $\mathbf{U}_{0}$ be $k\times d$ left singular vector matrix corresponding to the largest $d$ singular values in the singular value decomposition (SVD) of the Green’s function $\mathbf{G}\approx\mathbf{U}_{0}\mathbf{D}_{0}\mathbf{V}_{0}^{T}$, where $\mathbf{D}_{0}$ is a $d\times d$ diagonal matrix and $\mathbf{V}_{0}$ is a $k^{\prime}\times d$ matrix of $d$ right singular vectors with $d\leq k$. As the number of discretization points $k^{\prime}$ is larger than $k$, the latent slip vector is modeled by a linear combination of the latent processes [33] to obtain a reduced order representation:
(2) $\mathbf{z}_{s}(t)={\mathbf{G}}^{T}\mathbf{U}_{0}\mathbf{\tilde{z}}(t),$
@@
with $\mathbf{\tilde{z}}(t)$ being a $d$-vector of latent processes. The latent process $\tilde{z}_{l}(t)$ is assumed to follow an integrated Brownian motion in [33]. The Kalman filter is used for computing the posterior distribution of the latent process $\mathbf{\tilde{z}}(\cdot)$ and slip estimation. There are several restrictions in this approach. First, the latent process ${\tilde{z}}_{l}(\cdot)$ has the same covariance across $l=1,...,d$, yet the random latent factors can have different correlation length scales. Second, one needs to invert a $k\times k$ covariance matrix at each time in the Kalman filter, which is prohibitively slow when the number of GPS stations is moderately large. Third, selecting the number of latent factors is not discussed in this approach.
@@
The goal of this work is to develop a generally applicable approach and scalable algorithm for latent models for estimating the mean of the data and the posterior distribution of the latent variables. For estimating the slip rates in Cascadia between June and August in 2011, we found that the estimated slips had higher spatial correlation with independently detected tremor events not used in estimation. Furthermore, our method is $500-1600$ times faster than the current method in our real application. The high scalability of our method enables abundant data from GPS networks to be jointly used for hazard quantification.
@@ 2.2 A multivariate Ornstein-Uhlenbeck process with an orthogonal basis matrix
We follow the assumption in (2), which indicates that the mean in (1) is approximated by ${\mathbf{G}}\mathbf{z}_{s}(t)\approx\mathbf{U}_{0}\mathbf{z}(t)$, where $\mathbf{z}(t)=\mathbf{D}^{2}_{0}\mathbf{\tilde{z}}(t)$ with the $l$th entry being $z_{l}(t)$, and there is no approximation when $k=d$. We model the $d$-dimensional latent factor process $\mathbf{z}(\cdot)$ by independent Ornstein-Uhlenbeck (OU) processes, where each process has its own correlation and variance parameters. Putting together, Equations (1)-(2) lead to
(3) $\mathbf{y}(t)$ $=\mathbf{U}_{0}{\mathbf{z}}(t)+\bm{\epsilon}(t),$
(4) ${z}_{l}(t)$ $=\rho_{l}{z}_{l}(t-1)+w_{l}(t),$
where $\mathbf{U}_{0}$ is a $k\times d$ orthogonal factor loading matrix, either fixed or estimated, which relates the $d$-dimensional latent factors to $k$-dimensional measurements with $d\leq k$. Here, the parameter $\rho_{l}$ controls the temporal correlation length of the $l$th latent process, and $w_{l}(t)\sim\mathcal{N}(0,\sigma^{2}_{l})$ represents the innovation of the $l$th latent process for $t=2,...,n$, with the initial state being $z_{l}(1)\sim\mathcal{N}(0,\tau^{2}_{l})$ for $l=1,\dots,d$. The $d$ set of parameters $\bm{\rho}=(\rho_{1},\rho_{2},\dots,\rho_{d})$ and $\bm{\sigma}^{2}=(\sigma_{1}^{2},\sigma_{2}^{2},\dots,\sigma_{d}^{2})$ make the model flexible. We postpone the discussion of estimating $d$ in Section 2.4. Note that the matrix $\mathbf{G}$ in Equation (1) is typically neither symmetric nor orthogonal, whereas $\mathbf{U}_{0}$ is an orthogonal matrix, which brings substantial computational advantages, as will be elaborated soon. In this work, we develop novel algorithms for both fixed and estimated $\mathbf{U}_{0}$, which are broadly applicable to many applications. Furthermore, we assume the mean to be zero for simplicity, though additional mean structures can be included in the model.
Lemma 2.1 connects the model in (3)-(4) to the continuous-time multivariate Ornstein-Uhlenbeck process [12, 26] with an orthogonal basis matrix. The derivations of all lemmas in this subsection are provided in Section SM1 in supplementary materials.
@@
(5) $d\mathbf{m}(t)=-\mathbf{U}_{0}\mathbf{D}\mathbf{U}^{T}_{0}\mathbf{m}(t)dt+\mathbf{U}_{0}\mathbf{\tilde{D}}d\mathbf{B}_{t},$
@@
We study the model in (3)-(4) because of its flexibility due to distinct parameters $(\rho_{l},\sigma_{l}^{2})$ for each latent factor $l$, for $l=1,...,d$ and the computational scalability when the number of latent factors, $d$, is large. The latent process $z_{l}(\cdot)$ can be either stationary or nonstationary, depending on the assumption of the initial state, stated in the Lemma 7.
@@
(6) $\mbox{Cov}[z_{l}(t),z_{l}(t^{\prime})]=\rho_{l}^{t^{\prime}-t}\left\{\rho_{l}^{2(t-1)}\tau_{l}^{2}+\frac{1-\rho_{l}^{2(t-1)}}{(1-\rho_{l}^{2})}\sigma_{l}^{2}\right\},\quad\text{for }t^{\prime}\geq t.$
In particular, when the variance of initial state is $\tau^{2}_{l}=\frac{\sigma^{2}_{l}}{1-\rho^{2}_{l}}$, $z_{l}(\cdot)$ is stationary with
(7) $\mbox{Cov}[z_{l}(t),z_{l}(t^{\prime})]=\frac{\sigma^{2}_{l}}{1-\rho^{2}_{l}}\rho_{l}^{|t-t^{\prime}|}.$
In the rest of the article, we assume that each latent process is stationary by letting $\tau^{2}_{l}=\frac{\sigma^{2}_{l}}{1-\rho^{2}_{l}}$. Lemma 2.3 outlines key properties of the covariance and precision matrices, which enable us to modify Kalman filter (KF) [21] and Rauch–Tung–Striebel (RTS) smoother [30] to compute several key quantities in our fast algorithm.
@@
Assume $\tau^{2}_{l}=\frac{\sigma^{2}_{l}}{1-\rho^{2}_{l}}$ and denote $\mathbf{\Sigma}_{l}=\frac{\sigma_{l}^{2}}{1-\rho_{l}^{2}}\mathbf{R}_{l}$ as the covariance matrix of $\mathbf{z}_{l}=(z_{l}(1),\dots,z_{l}(n))^{T}$ defined in Equation (4). The $(t,t^{\prime})$ entry of $\mathbf{R}_{l}$ is $\rho_{l}^{|t-t^{\prime}|}$, for $t,t^{\prime}=1,2,\dots,n$. The inverse of $\mathbf{R}_{l}$ has a tri-diagonal structure, with diagonal entries of $\frac{1}{1-\rho_{l}^{2}}$ at the first and last positions, and $\frac{1+\rho_{l}^{2}}{1-\rho_{l}^{2}}$ at the remaining positions. The primary off-diagonal entries are $-\frac{\rho_{l}}{1-\rho_{l}^{2}}$. Additionally, the determinant of $\bm{\Sigma}_{l}$ follows $|\bm{\Sigma}_{l}|=\frac{\sigma^{2n}}{1-\rho_{l}^{2}}$.